A curtain is to be hung in a hospital room to give privacy. The path the curtain will follow is a rectangle with the wall being one side. If 27 ft of the curtain is used the the side parallel to the wall is 3 ft shorter than the two longer sides, find the dimensions of the curtained room.
Answer by Earlsdon(6294)

Let x = the longer side of the curtain enclosure, then:
x+x+(x-3) = 27 Simplify.
3x-3 = 27 Add 3 to both sides.
3x = 30 Divide by 3.
x = 10 and x-3 = 7.